Given
C= 60+0.75(1-0.2)y
Ī=280
G= 310
X= 250
Z= 150+0.25Y
A) calculate the equilibrium level of income in this open economy.
B) If the economy were closed(3 sector model), what would the value of the multiplier be?
C) if the current equilibrium level of income and spending is equal to 1615 million, what does the current tax income if the government? What type of government budget does the country have?
D) draw a graph for the tax function of the country. Describe the tax system in the economy. Show the slope of the graph.
1
Expert's answer
2020-06-01T12:24:54-0400

a)

Y= 60+0.75(1-0.2)y+280+310+250-(150+0.25Y)

Y=60+0.6y+280+310+250-150-0.25Y

Y-0.6y+0.25Y=750

0.65Y=750

Y=1 153.85


b)Y= 60+0.75(1-0.2)y+280+310

MPC=0.75

k= 1/(1 -МРС)=1/(1-0.75)=1/0.25=4


c)Y=1615

t=0.2

T=tY=0.2*1615=323

Saldo=T-G=323-310=13

budget surplus

d)





Proportional tax system. Revenue growth leads to higher taxes